# CATS: Enhancing Multivariate Time Series Forecasting by Constructing Auxiliary Time Series as Exogenous Variables

### v1

@@ Abstract
For Multivariate Time Series Forecasting (MTSF), recent deep learning applications show that univariate models frequently outperform multivariate ones. To address the difficiency in multivariate models, we introduce a method to Construct Auxiliary Time Series (CATS) that functions like a 2D temporal-contextual attention mechanism, which generates Auxiliary Time Series (ATS) from Original Time Series (OTS) to effectively represent and incorporate inter-series relationships for forecasting. Key principles of ATS—continuity, sparsity, and variability—are identified and implemented through different modules. Even with a basic 2-layer MLP as core predictor, CATS achieves state-of-the-art, significantly reducing complexity and parameters compared to previous multivariate models, marking it an efficient and transferable MTSF solution.
@@ 1 Introduction
Time series forecasting plays a pivotal role in domains like finance, health, and traffic, focusing on predicting future data based on historical records (Kaastra & Boyd, 1996; Nihan & Holmesland, 1980; Sun & Yang, 2023; Ma et al., 2022). The complexity increases in multivariate time series forecasting (MTSF), where understanding the interdependencies among multiple series is needed for accuracy. Effectively modeling these relationships is key to overcoming challenges such as overfitting, thus enhancing the reliability of predictions.
@@
Recent research in MTSF with deep learning methods reveals a counter-intuitive trend: univariate models, which ignore inter-series relationships, often outperform their multivariate counterparts (Zeng et al., 2023; Nie et al., 2022; Lu et al., 2023). This phenomenon suggests two potential issues: 1) The real-world benchmark datasets selected for study may inherently possess weak inter-series relationships; 2) Previous models may lack the proper structure needed to effectively capture inter-series relationships. This inadequacy often leads to overfitting or fitting of wrong patterns. Moreover, in situations where intra-series (temporal) relationships dominate, attempts to incorporate inter-series information may inadvertently disrupt the learning of temporal dynamics, resulting in the multivariate under-performance. This highlights the need for models that can balance and harness both types of relationships in MTSF.
@@
In this paper, we introduce an approach to effectively capture the two types of relationships in MTSF: Constructing Auxiliary Time Series (CATS). This approach empowers a time series predictor to handle diverse strengths of multivariate relationships, without altering its model architecture, even when the predictor is a structurally simple univariate model. CATS operates by generating Auxiliary Time Series (ATS) based on the Original Time Series (OTS), serving as exogenous variables to represent the inter-series relationships. By predicting these exogenous time series, CATS finds and incorporates multivariate information to the prediction of OTS, thereby boosting overall performance.
@@
We define ATS as the intermediary time series that are derived from the OTS input, reside in the same temporal domain as the original series. ATS can refine the univariate forecasting results with inter-series information. There are three key principles of ATS: continuity, sparsity, and variability. These principles are crucial when the inter-series relationships are weak, complex, or challenging to learn. ATS are constructed with these three principles using specific neural network structures, gating operations, and loss terms. This construction allows the ATS to effectively extract and represent multivariate relationships.
1.1 Basic Intuition: How CATS Works?
We present an intuitive example to demonstrate the efficacy and working mechanism of CATS. We apply the simplest form of ATS constructor – identity mapping – and use a basic independent linear predictor for prediction. Despite the ATS merely replicates the OTS, it remarkably transforms a univariate model into a model capable of capturing challenging multivariate relationships, through a mechanism similar to 2D temporal-contextual attention.
The example is a simple shifting problem introduced by (Lu et al., 2023) using a Multi dataset, where the best solution is a copy-paste operation. However, most multivariate models struggle due to challenges in timestep mixing as well as managing relationship across different time steps of multivariate series (see the Multi results in Table 1 and 7).
Figure 1: The Basic CATS Solution of the Shifting Problem
Using the CATS structure, as shown in Figure 1, we initially duplicate the first series to create an ATS, denoted as $A_{I}=X_{I}^{1}$. For the prediction, we use independent linear layers for each of $X^{1}$, $X^{2}$, and $A$. The best independent predictors for $X^{1}$ and $X^{2}$ will remain as their last values. However, the ATS’s linear layer operates differently. Let the weight matrix for this layer be denoted by $W\in\mathbb{R}^{48\times 144}$. When $W$ is structured such that its entries $W_{[:,49:96]}=I$ and 0 otherwise, this layer effectively ”transports” the steps 49 to 96 of the first series to the ATS’s output with $\widehat{A}_{Uni}=WA_{I}=X^{(49:96),1}_{I}$. Subsequently, a linear mapping then ”transports” this segment from the ATS’s output to the OTS output domain, making $\widehat{X}^{2}_{P}=X^{(49:96),1}_{I}$. In this simplified setting, even without complex network structure, CATS adaptively learns the inter-series relationship. This demonstrates CATS as a adaptive mechanism that focuses on specific regions of the OTS, summarize the inter-series information in ATS, and effectively transfers it from the input to the output domain, finally integrating it into the OTS output through projection.
Figure 2: 2D Temporal-contextual Attention-like Mechanism
The key contributions of this paper can be summrized as:
@@ 1.1 Basic Intuition: How CATS Works?
1) We introduce the Constructing ATS (CATS) approach based on the multivariate OTS, which empowers time series predictors with the ability of efficient capturing the inter-series relationships with a 2D attention-like mechanism.
2) We identify three key principles essential for the efficacy of ATS: continuity, sparsity, and variability, and design specific modules that enforce these principles.
@@ 2 Related Works
Time series forecasting models have been extensively studied in traditional statistical literature (Box et al., 1974; Holt, 2004; Sims, 1980). The advent of deep learning methods has provided time series forecasting with enhanced fitting capabilities, especially those based on RNN and CNN (Hochreiter & Schmidhuber, 1997; Wen et al., 2017; Rangapuram et al., 2018; Salinas et al., 2020; Qin et al., 2017; Borovykh et al., 2017; van den Oord et al., 2016), as well as more recent Transformer-based complex MTSF models (Wen et al., 2022). However, studies have pointed out that simpler univariate models, which ignore the relationships between series, can surpass these complex models (Zeng et al., 2023; Nie et al., 2022). This indicates that previous complex models may not effectively model these relationships.
For MTSF, effectively handling both temporal and contextual dimensions is crucial. Nonetheless, recent complex models are typically driven by 1D temporal attention mechanisms, as seen in LogTrans (Li et al., 2019), Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), etc, which has been shown to inadequately capture series-wise relationships (Zeng et al., 2023; Lu et al., 2023). Some research also tried to explicitly model series-wise relationships (Liu et al., 2023), but the input temporal embedding diminish the effectiveness with longer input length. Crossformer (Zhang & Yan, 2023) tried to build a 2D attention mechanism, but its performance did not surpass linear univariate models. ARM (Lu et al., 2023) attempted to simulate 2D attention by weighting the results of temporal attention with channel attention scores, but its complexity and reliance on univariate components limit its effectiveness.
@@ 3 Method
The CATS structure, as shown in Figure 3, while conducting the prediction of the original series, constructs ATS capable of representing inter-series relationships and incorporate the ATS predicting results into the OTS prediction to obtain the model output. In the absence of any ATS constructor, the output of CATS is identical to the result obtained by directly using a time series predictor. This predictor can be a univariate forecasting model or a multivariate forecasting model with neural network architecture. The time series predictor is denoted as $\Phi_{k}:\mathbb{R}^{L_{I}\times k}\rightarrow\mathbb{R}^{L_{P}\times k}$, which is capable of mapping a specified $k$ number of time series inputs from input to forecasting horizon. The prediction process of the model without ATS is expressed as $\widehat{X}_{P}=\Phi_{C}(X_{I})$. Let []delimited-[][$\ldots$][ … ] denotes concatenation along the series channel dimension. With ATS $A$, the predicting process for the combined set of $N+C$ time series becomes $[\widehat{A}_{P},\widehat{X}_{P}]=\Phi_{N+C}([A_{I},X_{I}])$, where
$A_{I}=[F_{1}(X_{I}),F_{2}(X_{I}),\ldots,F_{M}(X_{I})]\in\mathbb{R}^{L_{P}\times N}.$
@@
The original $\widehat{X}_{P}$ part of the prediction [^1] For distinction purposes, we use the term prediction to denote the computation results of ATS and OTS over $L_{P}$, and the term forecasting for the combined final output of the model. is retained as an OTS shortcut, encompassing most of the intra-series temporal relationship. For the remaining $\widehat{A}_{P}$, concatenated with $\widehat{X}_{P}$, we then project the $N+C$ result channels at each timestep into the $C$-dimensional space of the OTS, yielding the $\widetilde{X}_{P}$ that primarily captures the multivariate relationship. The process can be expressed as: $\widetilde{X}_{P}^{(t),i}=\sum_{j=1}^{N+C}P_{ij}\cdot[\widehat{A}_{P}^{(t)},\widehat{X}_{P}^{(t)}]^{j}$, where $\widetilde{X}_{P}$ denotes the integrated prediction, and $P$ denotes the linear mapping matrix.
Finally, we can directly add the inter-series relationship component as the residual part to merge the prediction results as $\widehat{X}^{*}_{P}=\widehat{X}_{P}+\widetilde{X}_{P}$, where $\widehat{X}^{*}_{P}$ represents the final forecasting after correcting with multivariate information.
@@ 3.1 Three Key Principles of ATS
This section discusses three critical principles—continuity, sparsity, and variability—that substantially boost model performance. These principles are critical for ATS to effectively extract useful information and accurately represent inter-series dependencies for correct pattern fitting.
@@ 3.1.1 Continuity
where $\beta_{cont}$ is a weighting factor that adjusts the influence of the continuity loss in the overall loss function. $\sigma_{n}$ is standard deviation of the ATS for channel $n$.
By applying continuity loss, ATS will smooth out temporal details and noise, focusing on overall trends that can be described by inter-series dependencies, leaving the remaining intra-series details to be better modeled by established methods on the OTS. This clear division of modeling focus using ATS and OTS enhances the reliability and reduces the complexity of forecasting. We further investigate this from two theoretical perspectives. Detailed explanation of this part can be found in Appendix B.1.
@@
Therefore, with continuity loss, the ATS prediction component $\widetilde{X}_{P}$ will primarily extract inter-series trends, allowing the OTS prediction component $\widehat{X}_{P}$ to focus on the more stable intra-series autocorrelation. This aligns perfectly with the structure and objectives of CATS.
@@ 3.1.2 Sparsity
Temporal Sparsity   Previous MTSF research has shown that altering the length $L_{I}$ significantly impacts model performance, leading most models to conduct hyper-parameter tuning of $L_{I}$ (Wu et al., 2021; Zhou et al., 2022; Zeng et al., 2023; Nie et al., 2022; Lu et al., 2023). Choosing a too short $L_{I}$ may fail to learn the correct dependencies, while a too long $L_{I}$ can result in overfitting. Clearly, a fixed $L_{I}$ for all the series is not the optimal choice. We need a mechanism to adaptively cutoff the time series input of the predictor. In this scenario, each channel can discard different lengths of long-term parts that are unnecessary for its prediction. Hence, the predictor input will include time series of varying lengths. We refer to this principle as temporal sparsity.
@@
$Q^{\prime(t)}_{i}=Q^{(t)}_{i}\cdot\left(h_{i}(t)-\mathtt{sg}\left(h_{i}(t)+\mathbf{1}_{\{h_{i}(t)>0\})}\right)\right)$
@@ 3.2 ATS Constructors
Convolution is a prevalent temporal feature extractor that is capable of learning the contribution of local patterns to forecasting. The impact of varying kernel sizes in convolution is well-documented. (Lai et al., 2018; Borovykh et al., 2017; van den Oord et al., 2016; Li et al., 2019; Lu et al., 2023). CATS employs convolution layers with different kernel sizes as ATS constructors to accommodate different scenarios. The $m$th constructor utilizes 1D convolution $\mathtt{Conv1d}_{[C\rightarrow n_{m}]}(\cdot,K,D,S,G)$, defined by kernel size $K$, padding $D$, stride $S$, channel groups $G$, and the $n_{m}$ output channels mapped from $C$ OTS. The constructor is given by: $F^{[\mathtt{Conv}]}_{m}(X_{I})=\mathtt{Conv1d}_{[C\rightarrow n_{m}]}(X_{I}^{\top},K,\lfloor\frac{K-1}{2}\rfloor,1,1)^{\top}$.
@@
Linear Projection This fundamental feature extractor linearly maps OTS across channels to form latent auxiliary series. Frequently used in MTSF models like Transformers and MLP-Mixers for input processing (Zhou et al., 2021; Wu et al., 2021; Zhang & Yan, 2023), linear projection offers a straightforward dimensional transformation. However, it can mix unrelated series, a risk CATS mitigates by combining various ATS constructors and retaining an OTS shortcut. This constructor is expressed as: $F^{[\mathtt{Lin}]}_{m}(X_{I})=\mathtt{Conv1d}_{[C\rightarrow n_{m}]}(X_{I}^{\top},1,0,0,1)^{\top}$

### v2

@@ Abstract
For Multivariate Time Series Forecasting (MTSF), recent deep learning applications show that univariate models frequently outperform multivariate ones. To address the deficiency in multivariate models, we introduce a method to Construct Auxiliary Time Series (CATS) that functions like a 2D temporal-contextual attention mechanism, which generates Auxiliary Time Series (ATS) from Original Time Series (OTS) to effectively represent and incorporate inter-series relationships for forecasting. Key principles of ATS—continuity, sparsity, and variability—are identified and implemented through different modules. Even with a basic 2-layer MLP as the core predictor, CATS achieves state-of-the-art, significantly reducing complexity and parameters compared to previous multivariate models, marking it as an efficient and transferable MTSF solution. The code implementation is available at this link
@@ 1 Introduction
Time series forecasting plays a pivotal role in domains like finance, health, and traffic, focusing on predicting future data based on historical records (Kaastra and Boyd, 1996; Nihan and Holmesland, 1980; Sun and Yang, 2023; Ma et al., 2022). The complexity of multivariate time series forecasting (MTSF) increases due to the need for understanding dependencies among multiple series. Effectively modeling these relationships is key to overcoming challenges such as overfitting, thus enhancing the reliability of predictions.
@@
Recent research in MTSF with deep learning methods reveals a counter-intuitive trend: univariate models, which ignore inter-series relationships, often outperform their multivariate counterparts (Zeng et al., 2023; Nie et al., 2022; Lu et al., 2024). This phenomenon suggests two potential issues: 1) The real-world benchmark datasets selected for study may inherently possess weak inter-series relationships; 2) Previous models may lack the proper structure needed to effectively capture inter-series relationships. This inadequacy often leads to overfitting or fitting of wrong patterns. Moreover, in situations where intra-series (temporal) relationships dominate, attempts to incorporate inter-series information may inadvertently disrupt the learning of temporal dynamics, resulting in multivariate underperformance. This highlights the need for models that can balance and harness both types of relationships in MTSF.
@@
In this paper, we introduce an approach to effectively capture the two types of relationships in MTSF: Constructing Auxiliary Time Series (CATS). This approach empowers a time series predictor to handle diverse strengths of multivariate relationships without altering its model architecture, even when the predictor is a structurally simple univariate model. CATS operates by generating Auxiliary Time Series (ATS) based on the Original Time Series (OTS), serving as exogenous variables to represent the inter-series relationships. By predicting these exogenous time series, CATS finds and incorporates multivariate information to the prediction of OTS, thereby boosting overall performance.
@@
We define ATS as the intermediary time series that are derived from the OTS input, reside in the same temporal domain as the original series. ATS can refine the univariate forecasting results by incorporating inter-series information. There are three key principles of ATS: continuity, sparsity, and variability. These principles are crucial when the inter-series relationships are weak, complex, or challenging to learn, and allow the model to adaptively adjust itself to suit datasets with varied characteristics. ATS are constructed with these three principles using specific neural network structures, gating operations, and loss terms. This construction allows the ATS to effectively extract and represent multivariate relationships.
The key contributions of this paper can be summarized as:
@@
1) We introduce the CATS approach, which constructs ATS based on the multivariate OTS and empowers time series predictors to efficiently capture inter-series relationships through a 2D attention-like mechanism.
2) We identify three key principles—continuity, sparsity, and variability—that enhance the performance and flexibility of ATS, enabling adaptive and accurate representation of inter-series dependencies and facilitating correct pattern fitting across various datasets.
@@ 2 Related Works
Time series forecasting models have been extensively studied in traditional statistical literature (Box et al., 1974; Holt, 2004; Sims, 1980). The advent of deep learning methods has provided time series forecasting with enhanced fitting capabilities, especially those based on RNN and CNN (Hochreiter and Schmidhuber, 1997; Wen et al., 2017; Rangapuram et al., 2018; Salinas et al., 2020; Qin et al., 2017; Borovykh et al., 2017; van den Oord et al., 2016), as well as more recent Transformer-based complex MTSF models (Wen et al., 2022). However, studies have pointed out that simpler univariate models, which ignore the relationships between series, can surpass these complex models (Zeng et al., 2023; Nie et al., 2022). This indicates that previous complex models may not effectively model these relationships.
For MTSF, effectively handling both temporal and contextual dimensions is crucial. Nonetheless, recent complex models are typically driven by 1D temporal attention mechanisms, as seen in LogTrans (Li et al., 2019), Informer (Zhou et al., 2021), Autoformer (Wu et al., 2021), etc, which has been shown to inadequately capture series-wise relationships (Zeng et al., 2023; Lu et al., 2024). Some research also tried to explicitly model series-wise relationships (Liu et al., 2023), but the input temporal embedding diminish the effectiveness with longer input length. Crossformer (Zhang and Yan, 2023) tried to build a 2D attention mechanism, but its performance did not surpass linear univariate models. ARM (Lu et al., 2024) attempted to simulate 2D attention by weighting the results of temporal attention with channel attention scores, but its complexity and reliance on univariate components limit its effectiveness.
@@ 3 Method
The CATS structure, as shown in Figure 1, while conducting the prediction of the original series, constructs ATS capable of representing inter-series relationships and incorporate the ATS first-stage predicting results into the OTS first-stage prediction to obtain the model output. In the absence of any ATS constructor, the output of CATS is identical to the result obtained by directly using a time series predictor. This predictor can be a univariate or multivariate forecasting model with neural network architecture. The time series predictor is denoted as $\Phi_{k}:\mathbb{R}^{L_{I}\times k}\rightarrow\mathbb{R}^{L_{P}\times k}$, which is capable of mapping a specified $k$ number of time series inputs from input to forecasting horizon. The prediction process of the model without ATS is expressed as $\widehat{X}_{P}=\Phi_{C}(X_{I})$. Let $\ldots$] denotes concatenation along the series channel dimension. With ATS $A$, the predicting process for the combined set of $N+C$ time series becomes $\widehat{A}_{P},\widehat{X}_{P}]=\Phi_{N+C}([A_{I},X_{I}])$, where $A_{I}=[F_{1}(X_{I}),F_{2}(X_{I}),\ldots,F_{M}(X_{I})]\in\mathbb{R}^{L_{P}\times N}.$
@@
The original $\widehat{X}_{P}$ part of the first-stage prediction is retained as an OTS shortcut, encompassing most of the intra-series temporal relationship. For the remaining $\widehat{A}_{P}$, concatenated with $\widehat{X}_{P}$, we then project the $N+C$ result channels at each timestep into the $C$-dimensional space of the OTS, yielding the $\widetilde{X}_{P}$ that primarily captures the multivariate relationship. The process can be expressed as: $\widetilde{X}_{P}^{(t),i}=\sum_{j=1}^{N+C}P_{ij}\cdot[\widehat{A}_{P}^{(t)},\widehat{X}_{P}^{(t)}]^{j}$, where $\widetilde{X}_{P}$ denotes the integrated prediction and $P$ denotes the linear mapping matrix. Finally, we can directly add the inter-series relationship component as the residual part to merge the first-stage prediction results as $\widehat{X}^{*}_{P}=\widehat{X}_{P}+\widetilde{X}_{P}$, where $\widehat{X}^{*}_{P}$ represents the final forecasting after correcting with multivariate information.
3.1 Basic Intuition: How CATS Works
We present an intuitive example to demonstrate the efficacy and working mechanism of CATS. We apply the simplest form of ATS constructor – identity mapping – and use a basic independent linear predictor for first-stage prediction. Despite this ATS constructor merely replicates the OTS without any trainable parameter, it still remarkably transforms a univariate model into a model capable of capturing challenging multivariate relationships, through the mechanism similar to 2D temporal-contextual attention provided by CATS.
The example is a simple shifting problem introduced by (Lu et al., 2024) using a Multi dataset, where the best solution is a copy-paste operation. However, most multivariate models struggle due to challenges in timestep mixing as well as managing relationship across different time steps of multivariate series (see the Multi results in Table 1 and 7).
Figure 2: The Basic CATS Solution of the Shifting Problem
Using the CATS structure, as shown in Figure 2, we initially duplicate the first series to create an ATS, denoted as $A_{I}=X_{I}^{1}$. For the first-stage prediction, we use independent linear layers for each of $X^{1}$, $X^{2}$, and $A$. The best independent predictors for $X^{1}$ and $X^{2}$ will remain as their last values. However, the ATS’s linear layer operates differently. Let the weight matrix for this layer be denoted by $W\in\mathbb{R}^{48\times 144}$. When $W$ is structured such that its entries $W_{[:,49:96]}=I$ and 0 otherwise, this layer effectively ”transports” the steps 49 to 96 of the first series to the ATS’s output with $\widehat{A}_{Uni}=WA_{I}=X^{(49:96),1}_{I}$. Subsequently, a linear mapping then ”transports” this segment from the ATS’s output to the OTS output domain, making $\widehat{X}^{2}_{P}=X^{(49:96),1}_{I}$. In this simplified setting, even without complex network structure, CATS adaptively learns the inter-series relationship. This demonstrates CATS as a adaptive mechanism that focuses on specific regions of the OTS, summarize the inter-series information in ATS, and effectively transfers it from the input to the output domain, finally integrating it into the OTS output through projection.
Figure 3: 2D Temporal-contextual Attention-like Mechanism
@@ 3.2 Three Key Principles of ATS
This section discusses three critical principles: continuity, sparsity, and variability. These principles substantially enhance both the performance and flexibility of the model, enabling ATS to adaptively and accurately extract useful information and represent inter-series dependencies. This facilitates correct pattern fitting across different datasets.
@@ 3.2.1 Continuity
where $\beta_{cont}$ is a weighting factor that adjusts the influence of the continuity loss in the overall loss function. $\sigma_{n}$ is standard deviation of the ATS for channel $n$. CATS is trained using a Mean Squared Error (MSE) loss combined with the continuity loss. By applying continuity loss, ATS will smooth out temporal details and noise, focusing on overall trends that can be described by inter-series dependencies, leaving the remaining intra-series details to be better modeled by established methods on the OTS. This clear division of modeling focus using ATS and OTS enhances the reliability and reduces the complexity of forecasting. We further investigate this from two theoretical perspectives. Detailed explanation of this part can be found in Appendix B.1.
@@
Therefore, with continuity loss, the ATS first-stage prediction component $\widetilde{X}_{P}$ will primarily extract inter-series trends, allowing the OTS prediction component $\widehat{X}_{P}$ to focus on the more stable intra-series autocorrelation. This aligns perfectly with the structure and objectives of CATS.
@@ 3.2.2 Sparsity
Temporal Sparsity Previous MTSF research has shown that altering the length $L_{I}$ significantly impacts model performance, leading most models to conduct hyper-parameter tuning of $L_{I}$ (Wu et al., 2021; Zhou et al., 2022; Zeng et al., 2023; Nie et al., 2022; Lu et al., 2024). Choosing a too short $L_{I}$ may fail to learn the correct dependencies, while a too long $L_{I}$ can result in overfitting. Clearly, a fixed $L_{I}$ for all the series is not the optimal choice. We need a mechanism to adaptively cutoff the time series input of the predictor. In this scenario, each channel can discard different lengths of long-term parts that are unnecessary for its prediction. Hence, the predictor input will include time series of varying lengths. We refer to this principle as temporal sparsity.
@@
$Q^{\prime(t)}_{i}=Q^{(t)}_{i}\cdot\left(h_{i}(t)-\mathtt{sg}\left(h_{i}(t)-\mathbf{1}_{\{h_{i}(t)>0\})}\right)\right)$
@@ 3.3 ATS Constructors
Convolution is a prevalent temporal feature extractor that is capable of learning the contribution of local patterns to forecasting. The impact of varying kernel sizes in convolution is well-documented. (Lai et al., 2018; Borovykh et al., 2017; van den Oord et al., 2016; Li et al., 2019; Lu et al., 2024). CATS employs convolution layers with different kernel sizes as ATS constructors to accommodate different scenarios. The $m$th constructor utilizes 1D convolution $\mathtt{Conv1d}_{[C\rightarrow n_{m}]}(\cdot,K,D,S,G)$, defined by kernel size $K$, padding $D$, stride $S$, channel groups $G$, and the $n_{m}$ output channels mapped from $C$ OTS. The constructor is given by: $F^{[\mathtt{Conv}]}_{m}(X_{I})=\mathtt{Conv1d}_{[C\rightarrow n_{m}]}(X_{I}^{\top},K,\lfloor\frac{K-1}{2}\rfloor,1,1)^{\top}$.
@@
Linear Projection This fundamental feature extractor linearly maps OTS across channels to form latent auxiliary series. Frequently used in MTSF models like Transformers and MLP-Mixers for input processing (Zhou et al., 2021; Wu et al., 2021; Zhang and Yan, 2023), linear projection offers a straightforward dimensional transformation. However, it can mix unrelated series, a risk CATS mitigates by combining various ATS constructors and retaining an OTS shortcut. This constructor is expressed as: $F^{[\mathtt{Lin}]}_{m}(X_{I})=\mathtt{Conv1d}_{[C\rightarrow n_{m}]}(X_{I}^{\top},1,0,0,1)^{\top}$